# A Step Toward World Models: A Survey on Robotic Manipulation

### v1

@@ IV-B1 Data Engine
Large-scale human teleoperation datasets have greatly advanced robot learning [16, 31, 32, 112, 44, 35]. However, collecting such data is labor-intensive and limits coverage across diverse environments and tasks. Vision-based world models, particularly video world models, offer an alternative by learning environment dynamics and generating synthetic data. These models can be broadly divided, according to whether they are conditioned on actions, into static video generation models [38], which predict general future scenes, and action-conditioned video generation models, which simulate how actions change the environment. Beyond data generation, video-based world models increasingly support diverse tasks such as planning, policy learning, and action prediction, which will be reflected in the following content.
@@
Specifically, Du et al. [59] target to enable visual world imagination, action planning and generating video demonstration for training by learning a text-conditioned video generation model. Wu et al. [6] train a large-scale video world model to generate accurate and realistic simulated experiences, enabling video prediction, visual planning, and policy training. Jang et al. [38] propose to leverage video world models [97] to generate robot video data. They first fine-tune video world models on a target robot to capture the embodiment-specific dynamics and kinematics and prompt the model with to initial frames and language instruction to generate corresponding data. Pseudo-action labels are generated by means of either a latent action model [113] or an inverse dynamics model (IDM) [114]. Lu et al. [34] leverage 3D-GS reconstruction with Diffusion Transformers (DiTs) to effectively model 3D dynamics, which can promote future scenes generation to support imitation learning and reinforcement learning. Ye et al. [115] synthesize data from diverse perspectives to introduce variations in texture, illumination, viewpoints, physical properties, task diversity, and interaction patterns. Their approach includes: (i) re-rendering real trajectories with diverse visual content, (ii) generating viewpoint-consistent multi-camera scenes with pose adjustments, and (iii) synthesizing embodied interaction sequences, such as converting first-person human videos into robot-centric demonstrations. To ensure realism and avoid hallucination artifacts, the authors further leverage a set of quality assessment metrics that evaluate geometric consistency [116], multiview consistency [116], text–scene alignment [117], and physical plausibility [117]. When constructing world models for training data generation, it is unrealistic to expect any training distribution to encompass all possible configurations of the world. To handle this, Barcellona et al. [33] construct a compositional world model to generate novel demonstration data for training by combining Gaussian Splatting [118] and physics simulators. Equivariant transformation is leveraged to augment data, which modify both observations and the corresponding action sequences to ensure semantical consistency.
Support reinforcement learning (RL) based Robotics. Wang et al. [119] present a video-based world model capable of predicting future visual observations conditioned on VLA-generated actions. A VLM-guided instant reflector serves as a reward function that quantifies task completion through the semantic alignment between the predicted trajectory and the textual instruction. Despite recent progress, existing methods continue to face challenges in generating diverse and counterfactual data that remain physically plausible, thereby limiting the quality and diversity of synthetic datasets [44].
@@ IV-B2 Evaluation
Traditionally, robot control policies have been developed and evaluated using handcrafted physics simulators [120, 121, 122]. However, such simulators rely on simplified or manually engineered dynamics, which struggle to capture complex real-world phenomena, particularly high-DoF interactions, deformable objects, and other non-rigid or contact-rich scenarios [123, 124, 125]. Consequently, the resulting discrepancies between simulated and real environments, commonly referred to as the sim-to-real gap, have significantly hindered the deployment and generalization of robotic policies in practice [126, 127]. To handle this, world models potentially emerge as a scalable, reproducible, and informative tool, which reduce reliance on trial-and-error in the real world. Compared to other filed such as autonomous driving [128] and navigation [129], simulated evaluation of robotic manipulation remains difficult because of the highly varied and dynamic interactions that arise between the agent and its environment. Li et al. [130] leverage a video generative world model [97] to produce videos based on action representations from a policy network. A success detector [131] is then used to evaluate task completion from the generated videos and corresponding text prompts. Quevedo et al. [74] evaluate robot polices by means of Monte Carlo rollouts in the world model and take a vision-language model, i.e., GPT-4o [132], as the reward model. He et al. [133] introduce a frame-level control and a motion-reinforced training to improve action-following ability and temporal, dynamic consistency, enhancing the dynamic prediction and action responsiveness of world simulator. More valuable transitions are discovered for policy learning. Zhu et al. [43] construct a frame-level, action-conditioned video world model based on a Diffusion Transformer, enabling scalable policy evaluation, planning, and future-scene generation. Liao et al. [57] take an action-conditioned video generator as the core to model the spatial, temporal, and semantic regularities of real-world interactions that are fundamental to robotic manipulation. The base world model can support future scene generation, action predictions, data engine and closed-loop policy evaluation. Wang et al. [73] promote the versatility of video world models for policy evaluation, visual simulation, synthetic data generation by perform training on heterogeneous actions data with a shared spatial-temporal transformer.
Escontrela et al. [134] train an autoregressive transformer-based video prediction model and use the next-token likelihoods of the frozen model as a general reward function across diverse tasks.
@@ V-A1 Training Data Scarcity
V-A1a Leveraging Pre-trained Models.
Given the limited availability of training data, many approaches leverage existing pre-trained models. For example, Xiang et al. [72] bypass the need for training from scratch by integrating a pre-trained LLM and a pre-trained video model, requiring only lightweight fine-tuning. Zhu et al. [43] initialize IRASim with the pre-trained weights of OpenSora [93] to expedite training. Similarly, Sudhakar et al. [37] leverage a pre-trained diffusion model, while Wang et al. [9] utilize Stable Video Diffusion, fine-tuned with robotic videos to adapt to the robotics domain. Song et al. [42] further exploit the world knowledge embedded in pre-trained autoregressive video generation models such as NOVA [135].
V-A1b Incorporating Auxiliary Data Sources.
Some works tackle the shortage of robot data by using other available sources, such as human manipulation datasets. For instance, Zhi et al. [69] use both human and robot manipulation videos for training. However, these datasets often contain cluttered backgrounds and similar-looking objects. To address this, they apply optical flow constraints to make the learned representation embodiment-agnostic. Sudhakar et al. [37] leverage an automatic hand segmentation method to obtain agent-agnostic data for robot learning. Others resort to more diverse data. For example, Yang et al. [19] leverage diverse kinds of data, including objects, scenes, actions, motions, language, and motor control, and convert all actions into a common format.
@@
Instead of relying on real-world data, Deng et al. [136] synthesize large-scale action data to train their model. To address the scarcity of 4D data, the Aether team [23] generate RGB-D synthetic videos and develop a robust camera-pose annotation pipeline to reconstruct full 4D dynamics. Similarly, Zhen et al. [24] build a 4D embodied video dataset that combines synthetic data with ground-truth depth, normal information and real-world data with estimated depth and normal maps obtained from off-the-shelf estimators.
@@ V-A2 Heterogeneous action data
World models should be able to understand different forms of actions and embodiments to ensure their real-world applications. A basic strategy is to utilize diverse datasets for training. However, the inherent cross-domain and cross-embodiment nature of datasets lead to heterogeneous actions data, including action spaces, action frequencies, and action horizon. For example, diverse embodiment (e.g., different degrees of freedom across robotic arms) and control interface (end effector (EEF) position for arms) would lead to actions of different forms. To handle this, Zheng et al. [137] learn to capture their shared structural features to obtain the generic atomic behaviors by means of vision language models. Similarly, Zheng et al. [73] lean a share latent space for actions by decoupling observation and actions. More strategies can borrow from relevant fields [138, 139, 140].
@@ V-A3 Action Label Missing
Action-labeled data, which are essential for learning action-conditioned future predictions [19], are particularly scarce in real-world settings.
@@
To address the lack of action-labeled data, self-supervised learning techniques have been explored [36, 63, 30, 64]. For instance, Finn et al. [36, 63] propose to learn pixel-level motion in a self-supervised manner, while Ebert et al. [30, 64] leverage image-to-image registration between consecutive video frames to capture dynamics without explicit action labels. However, goal image-based learning presents several drawbacks: such goals are inconvenient for humans to specify, may over-constrain the desired behavior (leading to sparse rewards), or under-specify task-relevant information for non-goal-reaching tasks.
@@
Another approach to handling missing action labels is to infer them directly from unlabeled videos. More specifically, Bruce et al. [20, 141] employ latent action autoencoders to extract latent actions in a self-supervised manner. In their studies, Bruce et al. [20] sample actions uniformly, while Gao et al. [141] introduce biased action sampling to encourage broader exploration and enable action reuse across contexts. Jiang et al. [38] extract pseudo-actions using either a latent action model [113] or an inverse dynamics model (IDM) [114]. Du et al. [59, 75, 26, 70] learn from unlabeled videos by training inverse dynamics models to infer actions or their embeddings. Ren et al. [75] further integrate an inverse dynamics module into a latent dynamics model to leverage rich temporal representations, improving the temporal consistency of predicted actions. Villar et al. [26] predict latent actions from object-centric representations.
@@
Some works aim to leverage pre-trained video models. For instance, Rigter et al. [67] adapt a pre-trained video diffusion model for action-conditioned world modeling by training a lightweight adapter, which is then fine-tuned on a small set of domain-specific, action-labeled videos. Black et al. [68] similarly employ a pre-trained image-editing diffusion model to support video-based world modeling. In addition, Zhu et al. [71] design a unified world model that integrates the action and video diffusion processes within a unified transformer architecture using separate diffusion timesteps. This can enable learning from action-free video data. Ko et al. [70] utilize optical flow extracted from videos, thereby circumventing the need for explicit action labels.
@@ V-B1 Inputs
Language. Task instructions are usually given in language. Many methods use pretrained models such as CLIP [22, 70, 142, 61], Phi [143, 42], or conditional VAEs [42] to extract semantic representations from the instructions.
Visual data. Similarly, visual inputs are often processed using pre-trained visual encoders. For example, Tian et al. [61] leverage pre-trained Vision Transformers (ViTs) [144] to process image observations. Wu et al. [6] employ a conditional VQGAN that encodes only task-relevant dynamic information, such as the position and pose of moving objects, to reduce temporal redundancy across frames. An autoregressive, GPT-like transformer is then used to generate the next tokens, which are decoded into future frames.
Action data. Actions are sometimes represented as integer values, which lack the contextual richness. This limitation can prevent world models from accurately capturing the intended meaning behind actions. To address this, He et al. [133] propose representing actions through language templates that explicitly encode their semantic meaning. In many cases, actions are instead expressed in natural language, as noted above. While this enables richer semantic representations, it also introduces challenges, such as instruction-following ambiguity, which are discussed in Section V-B2a.
Diverse data inputs. Robots need to gain a structured understanding of the world by jointly considering diverse sensory inputs. To achieve this, Song et al. [42] embed images and robot actions into a unified physical space, enabling the model to capture the sequential evolution of both the robot and its environment. Hong et al. [18] incorporate visual, auditory, tactile, and thermal modalities, projecting them into a shared feature space where a language model generates subsequent states and action tokens.
@@ V-B2 Challenges
Instructions convey task goals and can take various forms, including linguistic directives (natural language or structured text), visual cues (sketches, images, or demonstration videos), and others. Compared to image-based goals, textual descriptions provide a more abstract, compositional, and flexible way of specifying objectives, enabling better generalization, clearer intent communication, and more efficient human–robot interaction. Many recent works express target goals through text descriptions [59]. Ideally, language instructions should clearly describe the task and remain easily interpretable by the model. However, real-world scenarios often involve ambiguous or novel instructions, making effective interpretation and grounding critical for successful task execution.
Ambiguous instructions In real-world scenarios, language instructions are often ambiguous (e.g., “put this near here” [9]). To resolve such ambiguity, Wang et al. [9] use pointing gestures, interpreted through 2D gripper and object tracking, as an additional instruction modality.
New instructions World models are constrained to make predictions based on language instructions similar to those encountered during training, limiting their ability to generalize to novel commands. To solve this problem, Xiang et al. [72] curate a large and diverse set of action-state sequences from re-captioned videos and simulations, and fine-tune world models on this data to improve instruction interpretation and generalize to novel commands and tasks. Li et al. [66] employ a text parser to decompose language instructions into primitives, separating actions and spatial relationships. This decomposition allows the model to flexibly recombine these components and generalize to previously unseen combinations of instructions. However, decomposing instructions into primitives can ignore their interrelationships. To address this, Li et al. [145] represent each instruction as an action tree, capturing the hierarchical structure among primitives to better model task organization.
@@
Some studies suggest that humans make predictions based on abstract concepts rather than raw pixels [27]. Instead of converting images into discrete tokens [19, 6], Chen et al. [27] use learnable convolutional layers to project images into continuous semantic embeddings. Song et al. [42] adopt an open-source 3D variational autoencoder (Open-Sora [93]) to obtain video representations. In contrast, another line of work operates directly in pixel space. For instance, Ko et al. [70] adapt a U-Net-based image diffusion model with factorized spatial–temporal convolutions [146] to jointly capture spatial and temporal information.
@@
Visual data often contain information irrelevant to the task, and models such as Vision Transformers (ViTs) may produce hundreds of features per image, affecting both efficiency and effectiveness. To address this, Tian et al. [61] extract task-relevant features using a perceiver resampler [147]. Ren et al. [75] learn compact visual representations that preserve fine-grained temporal dynamics through a causal encoder–decoder structure and quantization with a discrete codebook [148].
@@
Understanding the world requires modeling how spatial structures evolve over time. To this end, several works design architectures that explicitly capture spatial and temporal dependencies. Tian et al. [61] enhance token representations with learnable positional embeddings at each timestep to capture temporal information. Bruce et al. [20] develop a spatiotemporal transformer composed of multiple spatiotemporal blocks to model spatial–temporal relationships in dynamic scenes. Ko et al. [70] adopt factorized spatiotemporal convolutions following the design of [95]. Zhang et al. [60] extract spatiotemporal patch representations using a masked autoencoder [144]. Other studies incorporate additional cues to better understand the three-dimensional structure of the environment. For example, Zhang et al. [60] estimate depth information using depth estimation techniques [149] to enhance 3D spatial understanding. When encoding multi-view inputs, Liao et al. [57] augment each token with 2D rotary positional embeddings, view-specific learnable embeddings, and timestep encodings to promote spatiotemporal alignment while preserving viewpoint-specific distinctions.

### v2

@@ IV-B1 Data Engine
Large-scale human teleoperation datasets have greatly advanced robot learning [zitkovich2023rt, black2024pi_0, team2025gemini, bu2025agibot, bjorck2025gr00t, liu2025rdt]. However, collecting such data is labor-intensive and limits coverage across diverse environments and tasks. Vision-based world models, particularly video world models, offer an alternative by learning environment dynamics and generating synthetic data. These models can be broadly divided, according to whether they are conditioned on actions, into static video generation models [jang2025dreamgen], which predict general future scenes, and action-conditioned video generation models, which simulate how actions change the environment. Beyond data generation, video-based world models increasingly support diverse tasks such as planning, policy learning, and action prediction, which will be reflected in the following content.
@@
Specifically, Du et al. [du2023learning] target to enable visual world imagination, action planning and generating video demonstration for training by learning a text-conditioned video generation model. Wu et al. [wu2024ivideogpt] train a large-scale video world model to generate accurate and realistic simulated experiences, enabling video prediction, visual planning, and policy training. Jang et al. [jang2025dreamgen] propose to leverage video world models [wan2025wan] to generate robot video data. They first fine-tune video world models on a target robot to capture the embodiment-specific dynamics and kinematics and prompt the model with to initial frames and language instruction to generate corresponding data. Pseudo-action labels are generated by means of either a latent action model [ye2025latent] or an inverse dynamics model [baker2022video]. Lu et al. [lu2025gwm] leverage 3D-GS reconstruction with Diffusion Transformers to effectively model 3D dynamics, which can promote future scenes generation to support imitation learning and reinforcement learning. Ye et al. [Ye2025GigaBrain] synthesize data from diverse perspectives to introduce variations in texture, illumination, viewpoints, physical properties, task diversity, and interaction patterns. Their approach includes: (i) re-rendering real trajectories with diverse visual content, (ii) generating viewpoint-consistent multi-camera scenes with pose adjustments, and (iii) synthesizing embodied interaction sequences, such as converting first-person human videos into robot-centric demonstrations. To ensure realism and avoid hallucination artifacts, the authors further leverage a set of quality assessment metrics that evaluate geometric consistency [liu2025robotransfer], multiview consistency [liu2025robotransfer], text–scene alignment [azzolini2025cosmos], and physical plausibility [azzolini2025cosmos]. When constructing world models for training data generation, it is unrealistic to expect any training distribution to encompass all possible configurations of the world. To handle this, Barcellona et al. [barcellona2025dream] construct a compositional world model to generate novel demonstration data for training by combining Gaussian Splatting [kerbl20233d] and physics simulators. Equivariant transformation is leveraged to augment data, which modify both observations and the corresponding action sequences to ensure semantical consistency. Wang et al. [xiao2025world] present a video-based world model capable of predicting future visual observations conditioned on VLA-generated actions. A VLM-guided instant reflector serves as a reward function that quantifies task completion through the semantic alignment between the predicted trajectory and the textual instruction.
Despite recent progress, existing methods continue to face challenges in generating diverse and counterfactual data that remain physically plausible, thereby limiting the quality and diversity of synthetic datasets [bjorck2025gr00t].
@@ IV-B2 Evaluation
Traditionally, robot control policies have been developed and evaluated using handcrafted physics simulators [todorov2012mujoco, erez2015simulation, tedrake2019drake]. However, such simulators rely on simplified or manually engineered dynamics, which struggle to capture complex real-world phenomena, particularly high-DoF interactions, deformable objects, and other non-rigid or contact-rich scenarios [sunderhauf2018limits, afzal2020study, choi2021use]. Consequently, the resulting discrepancies between simulated and real environments, commonly referred to as the sim-to-real gap, have significantly hindered the deployment and generalization of robotic policies in practice [dulac2019challenges, zhao2020sim]. To handle this, world models potentially emerge as a scalable, reproducible, and informative tool, which reduce reliance on trial-and-error in the real world. Compared to other filed such as autonomous driving [dosovitskiy2017carla] and navigation [deitke2020robothor], simulated evaluation of robotic manipulation remains difficult because of the highly varied and dynamic interactions that arise between the agent and its environment. Li et al. [li2025worldeval] leverage a video generative world model [wan2025wan] to produce videos based on action representations from a policy network. A success detector [team2023gemini] is then used to evaluate task completion from the generated videos and corresponding text prompts. Quevedo et al. [quevedo2025evaluating] evaluate robot polices by means of Monte Carlo rollouts in the world model and take a vision-language model, i.e., GPT-4o [hurst2024gpt], as the reward model. He et al. [he2025pre] introduce a frame-level control and a motion-reinforced training to improve action-following ability and temporal, dynamic consistency, enhancing the dynamic prediction and action responsiveness of world simulator. More valuable transitions are discovered for policy learning. Zhu et al. [zhu2025irasim] construct a frame-level, action-conditioned video world model based on a Diffusion Transformer, enabling scalable policy evaluation, planning, and future-scene generation. Liao et al. [liao2025genie] take an action-conditioned video generator as the core to model the spatial, temporal, and semantic regularities of real-world interactions that are fundamental to robotic manipulation. The base world model can support future scene generation, action predictions, data engine and closed-loop policy evaluation. Wang et al. [wang2025learning] promote the versatility of video world models for policy evaluation, visual simulation, synthetic data generation by perform training on heterogeneous actions data with a shared spatial-temporal transformer.
Escontrela et al. [escontrela2023video] train an autoregressive transformer-based video prediction model and use the next-token likelihoods of the frozen model as a general reward function across diverse tasks.
@@ V-A1 Training Data Scarcity
V-A1a Leveraging Pre-trained Models
Given the limited availability of training data, many approaches leverage existing pre-trained models. For example, Xiang et al. [xiang2024pandora] bypass the need for training from scratch by integrating a pre-trained LLM and a pre-trained video model, requiring only lightweight fine-tuning. Zhu et al. [zhu2025irasim] initialize IRASim with the pre-trained weights of OpenSora [zheng2024open] to expedite training. Similarly, Sudhakar et al. [sudhakar2024controlling] leverage a pre-trained diffusion model, while Wang et al. [wang2025language] utilize Stable Video Diffusion, fine-tuned with robotic videos to adapt to the robotics domain. Song et al. [song2025physical] further exploit the world knowledge embedded in pre-trained autoregressive video generation models such as NOVA [deng2025autoregressive].
V-A1b Incorporating Auxiliary Data Sources
Some works tackle the shortage of robot data by using other available sources, such as human manipulation datasets. For instance, Zhi et al. [zhi20253dflowaction] use both human and robot manipulation videos for training. However, these datasets often contain cluttered backgrounds and similar-looking objects. To address this, they apply optical flow constraints to make the learned representation embodiment-agnostic. Sudhakar et al. [sudhakar2024controlling] leverage an automatic hand segmentation method to obtain agent-agnostic data for robot learning. Others resort to more diverse data. For example, Yang et al. [yang2023learning] leverage diverse kinds of data, including objects, scenes, actions, motions, language, and motor control, and convert all actions into a common format.
@@
Instead of relying on real-world data, Deng et al. [deng2025graspvla] synthesize large-scale action data to train their model. To address the scarcity of 4D data, the Aether team [team2025aether] generate RGB-D synthetic videos and develop a robust camera-pose annotation pipeline to reconstruct full 4D dynamics. Similarly, Zhen et al. [zhen2025tesseract] build a 4D embodied video dataset that combines synthetic data with ground-truth depth, normal information and real-world data with estimated depth and normal maps obtained from off-the-shelf estimators.
@@ V-A2 Heterogeneous Action Data
World models should be able to understand different forms of actions and embodiments to ensure their real-world applications. A basic strategy is to utilize diverse datasets for training. However, the inherent cross-domain and cross-embodiment nature of datasets lead to heterogeneous actions data, including action spaces, action frequencies, and action horizon. For example, diverse embodiment (e.g., different degrees of freedom across robotic arms) and control interface (end effector (EEF) position for arms) would lead to actions of different forms. To handle this, Zheng et al. [zheng2025universal] learn to capture their shared structural features to obtain the generic atomic behaviors by means of vision language models. Similarly, Zheng et al. [wang2025learning] lean a share latent space for actions by decoupling observation and actions. More strategies can borrow from relevant fields [doshi2025scaling, team2024octo, wang2024scaling].
@@ V-A3 Action Label Missing
Action-labeled data, which are essential for learning action-conditioned future predictions [yang2023learning], are particularly scarce in real-world settings.
@@
Finn et al. [finn2016unsupervised, finn2017deep] propose to learn pixel-level motion in a self-supervised manner, while Ebert et al. [ebert2018robustness, ebert2018visual] leverage image-to-image registration between consecutive video frames to capture dynamics without explicit action labels. However, goal image-based learning presents several drawbacks: such goals are inconvenient for humans to specify, may over-constrain the desired behavior (leading to sparse rewards), or under-specify task-relevant information for non-goal-reaching tasks.
@@
Another approach to handling missing action labels is to infer them directly from unlabeled videos. More specifically, Bruce et al. [bruce2024genie, gao2025adaworld] employ latent action autoencoders to extract latent actions in a self-supervised manner. In their studies, Bruce et al. [bruce2024genie] sample actions uniformly, while Gao et al. [gao2025adaworld] introduce biased action sampling to encourage broader exploration and enable action reuse across contexts. Jiang et al. [jang2025dreamgen] extract pseudo-actions using either a latent action model [ye2025latent] or an inverse dynamics model (IDM) [baker2022video]. Du et al. [du2023learning, ren2025videoworld, villar2025playslot, ko2024learning] learn from unlabeled videos by training inverse dynamics models to infer actions or their embeddings. Ren et al. [ren2025videoworld] further integrate an inverse dynamics module into a latent dynamics model to leverage rich temporal representations, improving the temporal consistency of predicted actions. Villar et al. [villar2025playslot] predict latent actions from object-centric representations.
@@
Some works aim to leverage pre-trained video models. For instance, Rigter et al. [rigter2025avid] adapt a pre-trained video diffusion model for action-conditioned world modeling by training a lightweight adapter, which is then fine-tuned on a small set of domain-specific, action-labeled videos. Black et al. [black2024zero] similarly employ a pre-trained image-editing diffusion model to support video-based world modeling. In addition, Zhu et al. [zhu2025unified] design a unified world model that integrates the action and video diffusion processes within a unified transformer architecture using separate diffusion timesteps. This can enable learning from action-free video data. Ko et al. [ko2024learning] utilize optical flow extracted from videos, thereby circumventing the need for explicit action labels.
Implications for Data Limitations The scarcity of data can be alleviated by leveraging external datasets or pre-trained base models from other sources. The key challenge lies in bridging the gap between the source and target domains. Furthermore, it is desirable to uncover the underlying knowledge through supervised learning.
@@ V-B1 Inputs
Language. Task instructions are usually given in language. Many methods use pretrained models such as CLIP [bu2024closed, ko2024learning, radford2021learning, tian2025predictive], Phi [javaheripi2023phi, song2025physical], or conditional VAEs [song2025physical] to extract semantic representations from the instructions.
Visual data. Similarly, visual inputs are often processed using pre-trained visual encoders. For example, Tian et al. [tian2025predictive] leverage pre-trained Vision Transformers (ViTs) [he2022masked] to process image observations. Wu et al. [wu2024ivideogpt] employ a conditional VQGAN that encodes only task-relevant dynamic information, such as the position and pose of moving objects, to reduce temporal redundancy across frames. An autoregressive, GPT-like transformer is then used to generate the next tokens, which are decoded into future frames.
Action data. Actions are sometimes represented as integer values, which lack the contextual richness. This limitation can prevent world models from accurately capturing the intended meaning behind actions. To address this, He et al. [he2025pre] propose representing actions through language templates that explicitly encode their semantic meaning. In many cases, actions are instead expressed in natural language, as noted above. While this enables richer semantic representations, it also introduces challenges, such as instruction-following ambiguity, which are discussed in Section V-B2a.
Diverse data inputs. Robots need to gain a structured understanding of the world by jointly considering diverse sensory inputs. To achieve this, Song et al. [song2025physical] embed images and robot actions into a unified physical space, enabling the model to capture the sequential evolution of both the robot and its environment. Hong et al. [hong2024multiply] incorporate visual, auditory, tactile, and thermal modalities, projecting them into a shared feature space where a language model generates subsequent states and action tokens.
@@ V-B2 Challenges
Instructions convey task goals and can take various forms, including linguistic directives (natural language or structured text), visual cues (sketches, images, or demonstration videos), and others. Compared to image-based goals, textual descriptions provide a more abstract, compositional, and flexible way of specifying objectives, enabling better generalization, clearer intent communication, and more efficient human–robot interaction. Many recent works express target goals through text descriptions [du2023learning]. Ideally, language instructions should clearly describe the task and remain easily interpretable by the model. However, real-world scenarios often involve ambiguous or novel instructions, making effective interpretation and grounding critical for successful task execution.
Ambiguous Instructions. In real-world scenarios, language instructions are often ambiguous (e.g., “put this near here” [wang2025language]). To resolve such ambiguity, Wang et al. [wang2025language] use pointing gestures, interpreted through 2D gripper and object tracking, as an additional instruction modality.
New Instructions. World models are constrained to make predictions based on language instructions similar to those encountered during training, limiting their ability to generalize to novel commands. To solve this problem, Xiang et al. [xiang2024pandora] curate a large and diverse set of action-state sequences from re-captioned videos and simulations, and fine-tune world models on this data to improve instruction interpretation and generalize to novel commands and tasks. Li et al. [zhou2024robodreamer] employ a text parser to decompose language instructions into primitives, separating actions and spatial relationships. This decomposition allows the model to flexibly recombine these components and generalize to previously unseen combinations of instructions. However, decomposing instructions into primitives can ignore their interrelationships. To address this, Li et al. [li2025manipdreamer] represent each instruction as an action tree, capturing the hierarchical structure among primitives to better model task organization.
@@
Some studies suggest that humans make predictions based on abstract concepts rather than raw pixels [chen2025egoagent]. Instead of converting images into discrete tokens [yang2023learning, wu2024ivideogpt], Chen et al. [chen2025egoagent] use learnable convolutional layers to project images into continuous semantic embeddings. Song et al. [song2025physical] adopt an open-source 3D variational autoencoder (Open-Sora [zheng2024open]) to obtain video representations. In contrast, another line of work operates directly in pixel space. For instance, Ko et al. [ko2024learning] adapt a U-Net-based image diffusion model with factorized spatial–temporal convolutions [dhariwal2021diffusion] to jointly capture spatial and temporal information.
@@
Visual data often contain information irrelevant to the task, and models such as Vision Transformers (ViTs) may produce hundreds of features per image, affecting both efficiency and effectiveness. To address this, Tian et al. [tian2025predictive] extract task-relevant features using a perceiver resampler [alayrac2022flamingo]. Ren et al. [ren2025videoworld] learn compact visual representations that preserve fine-grained temporal dynamics through a causal encoder–decoder structure and quantization with a discrete codebook [mentzer2024finite].
@@
Understanding the world requires modeling how spatial structures evolve over time. To this end, several works design architectures that explicitly capture spatial and temporal dependencies. Tian et al. [tian2025predictive] enhance token representations with learnable positional embeddings at each timestep to capture temporal information. Bruce et al. [bruce2024genie] develop a spatiotemporal transformer composed of multiple spatiotemporal blocks to model spatial–temporal relationships in dynamic scenes. Ko et al. [ko2024learning] adopt factorized spatiotemporal convolutions following the design of [ho2022video]. Zhang et al. [zhang2025dreamvla] extract spatiotemporal patch representations using a masked autoencoder [he2022masked]. Other studies incorporate additional cues to better understand the three-dimensional structure of the environment. For example, Zhang et al. [zhang2025dreamvla] estimate depth information using depth estimation techniques [yang2024depth] to enhance 3D spatial understanding. When encoding multi-view inputs, Liao et al. [liao2025genie] augment each token with 2D rotary positional embeddings, view-specific learnable embeddings, and timestep encodings to promote spatiotemporal alignment while preserving viewpoint-specific distinctions.